# Bidiagonal matrix factorisations related to multiple orthogonal polynomials

### v1

@@ 2. $\bullet$
$\alpha_{(r+1)(rm+k)+i}=\frac{\displaystyle\left(a_{k+i-r+1}-a_{i}+m+1\right)\prod_{j=0}^{i-1}\left(a_{j}+b+rm+k+2\right)\prod_{j=i+1}^{r}\left(a_{j}+b+rm+k+1\right)}{\displaystyle\prod_{j=1}^{k+i+1-r}\Big{(}a_{j}+b+(r+1)m+k+3\Big{)}\prod_{j=k+i+1-r}^{r}\Big{(}a_{j}+b+(r+1)m+k+2\Big{)}}\text{ if }k+i\geq r.$ (5.11)
Here, we find again the bidiagonal factorisation described above for the Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials, using explicit expressions for the type I polynomials recently given in [7]. It follows from [7, Thm. 2] that the leading term of the type I linear form $q_{rm+k-1}^{[i]}$, with $m\in\mathbb{N}$ and $1\leq k\leq r$, on the step-line with respect to the system $V^{[i]}$, $0\leq i\leq r$, is
$A^{[i],m}_{rm+k-1,k}=\frac{\displaystyle\Gamma\left(a^{*}_{k+i}+b+(r+1)m+k\right)\prod_{j=1}^{k}\left(a^{*}_{j+i}+b+rm+k\right)_{m+1}\prod_{j=k+1}^{r}\left(a^{*}_{j+i}+b+rm+k\right)_{m}}{\displaystyle m!\,\Gamma\left(a^{*}_{k+i}+m+1\right)\Gamma(b+rm+k)\prod_{j=1}^{k-1}\left(a^{*}_{k+i}-a^{*}_{j+i}\right)_{m+1}\prod_{q=k+1}^{r}\left(a^{*}_{k+i}-a^{*}_{j+i}+1\right)_{m}},$ (5.12)
where $a^{*}_{rn+j}=a_{j}+n$ for all $0\leq j\leq r$ and $n\in\mathbb{N}$, as in (5.4).
Inputting these coefficients in (3.38)-(3.39), we derive the formulas
$\alpha_{(r+1)(rm+k)}=\frac{\displaystyle(a^{*}_{k+1}+m+1)\prod\limits_{j=1}^{r}\left(a^{*}_{j}+b+rm+k+1\right)}{\displaystyle\prod\limits_{j=1}^{r+1}\left(a^{*}_{k+j}+b+(r+1)m+k+1\right)},$ (5.13)
and
$\alpha_{(r+1)(rm+k)+i}=\frac{\displaystyle(a^{*}_{k+i+1}-a^{*}_{i}+m)(b+rm+k+1)\prod\limits_{j=1}^{r-1}\left(a^{*}_{i+j}+b+rm+k+1\right)}{\displaystyle\prod\limits_{j=1}^{r+1}\left(a^{*}_{k+i+j}+b+(r+1)m+k+1\right)},$ (5.14)
for the nontrivial entries of the bidiagonal matrices appearing in the factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ of the Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials on the step-line. It is a straightforward computation to check that (5.10)-(5.11) and (5.13)-(5.14) are equivalent.
When $r=2$, the coefficients in (5.10)-(5.11) and (5.13)-(5.14) coincide with the ones found in [3, § 8.1].
It is a direct consequence of (5.10)-(5.11) or (5.13)-(5.14) that $\alpha_{n}\to\frac{r^{r}}{(r+1)^{r+1}}$ as $n\to\infty$. Therefore, recalling (2.10), the coefficients of the recurrence relation (2.7) satisfied by the type II Jacobi-Piñeiro polynomials on the step-line have the same limit as in [17, Eq. 8.17]: $\gamma_{n}^{[k]}\to\binom{r+1}{k+1}\left(\frac{r^{r}}{(r+1)^{r+1}}\right)^{k+1}$ as $n\to\infty$.
@@ 5.3. Multiple Laguerre of first kind as a limiting case
For a positive integer $r$ and parameters $a_{1},\cdots,a_{r}>-1$ such that $a_{i}-a_{j}\not\in\mathbb{Z}$ whenever $i\neq j$, we consider the measures $\left(\nu_{1},\cdots,\nu_{r}\right)$ supported on the interval $(0,+\infty)$ with densities $\mathrm{d}\nu_{j}(x)=\mathrm{e}^{-x}\,x^{a_{j}}\mathrm{d}x$, whose moments are
$\int_{0}^{+\infty}x^{n}\mathrm{d}\nu_{j}(x)=\int_{0}^{+\infty}\mathrm{e}^{-x}\,x^{a_{j}+n}\,\mathrm{d}x=\Gamma(a_{j}+n+1)\quad\text{for all }n\in\mathbb{N}\text{ and }1\leq j\leq r.$ (5.15)
@@
The Hessenberg recurrence matrix of the multiple Laguerre polynomials of the first kind on the step-line, multiple orthogonal with respect to $(\nu_{1},\cdots,\nu_{r})$ given by (5.15), admits the factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$, as a product of the bidiagonal matrices in (1.2) with nontrivial entries $\left(\hat{\alpha}_{n}\right)_{n\in\mathbb{N}}$ such that $\hat{\alpha}_{n}=\lim\limits_{b\to\infty}b\,\alpha_{n}$ for all $n\in\mathbb{N}$, where $\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ are the nontrivial entries in the bidiagonal decomposition of the recurrence matrix of the Jacobi-Piñeiro polynomials given in (5.10)-(5.11) and (5.13)-(5.14). Therefore, for any $m\in\mathbb{N}$, $10\leq k\leq r-$ 1 and $0\leq i\leq r$,
$\hat{\alpha}_{(r+1)(rm+k)+i}=a^{*}_{k+i+1}-a^{*}_{i}+m=\begin{cases}\hfil a_{k+i+1}-a_{i}+m&\text{if }k+i\leq r,\\ a_{k+i+1-r}-a_{i}+m&\text{if }k+i\geq r+1,\end{cases}$ (5.19)
with $a^{*}_{0}=a_{0}=-1$ and $a^{*}_{r+j}=a^{*}_{j}+1$ for any $1\leq j\leq r$.
As a consequence of (5.19), we know that $\hat{\alpha}_{n}\sim\frac{n}{r(r+1)}$ as $n\to\infty$. Therefore, recalling (2.10), the coefficients of the recurrence relation (2.7) satisfied by the type II multiple Laguerre polynomials of the first kind on the step-line satisfy the asymptotic behaviour $\hat{\gamma}_{n}^{[k]}\sim\binom{r+1}{k+1}\left(\frac{n}{r}\right)^{k+1}$ as $n\to\infty$.
Note that, alternatively to taking the limit in (5.19), we could try to apply the algorithms developed in Sections 3 and 4 directly to the multiple Laguerre polynomials of first kind. For the algorithm from Section 3, this process works without any issue: the formulas for the type I multiple Laguerre polynomials of the first kind with respect to an arbitrary number of weights can be found in [7, Thm. 3] and could be applied to (3.38)-(3.39) to obtain the coefficients $\left(\hat{\alpha}_{n}\right)_{n\in\mathbb{N}}$. However, it is much easier to simply take the limit in (5.19).
For the algorithm from Section 3, the situation is more complicated. In that case, the nontrivial entries $\left(\hat{\alpha}_{n}\right)_{n\in\mathbb{N}}$ of the bidiagonal decomposition of the recurrence matrix of the multiple Laguerre polynomials of first kind appear as the coefficients of the recurrence relation
@@
This relation is satisfied by the sequence $\left(\hat{g}_{k}(t)\right)_{k\in\mathbb{N}}$ obtained by taking the limit
$\hat{g}_{k}(t)=\lim_{b\to\infty}g_{k}(bt)=\lim_{b\to\infty}(1-bt)^{b}{}_{r+1}F_{r}\left(\left.\begin{matrix}a^{\,\prime}_{k}+b+1,\cdots,a^{\,\prime}_{k+r}+b+1\vspace*{2 pt}\\ a^{\,*}_{k+1}+b+1,\cdots,a^{\,*}_{k+r}+b+1\end{matrix}\,\right|\,bt\right).$ (5.22)
However, we could not calculate this limit and we would be interested in finding how to calculate it.
Conclusions and future work
In this work, we succeeded to develop two methods to obtain a bidiagonal factorisation for the Hessenberg recurrence matrix of a system of multiple orthogonal polynomials. One method is based on the Gauss-Borel factorisation of the moment matrix of the system and of its Darboux transformations and the nontrivial entries of the bidiagonal matrices are expressed as coefficients of the type I and type II multiple orthogonal polynomials on the step-line, while the other uses the connection of multiple orthogonal polynomials with production matrices and branched continued fractions and the nontrivial entries of the bidiagonal matrices are expressed as coefficients of the corresponding branched continued fraction. Furthermore, we used both methods to present an explicit bidiagonal factorisation for the Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials and, as a limiting case, the multiple Laguerre polynomials of first kind.
In forthcoming work, already initiated in [5], we plan to further extend the research presented here to the context of mixed multiple orthogonality.
@@ Acknowledgements
AB acknowledges Centre for Mathematics of the University of Coimbra funded by the Portuguese Government through FCT/MCTES, DOI: 10.54499/UIDB/00324/2020. JEFD, AF, and HL acknowledge Center for Research and Development in Mathematics and Applications (CIDMA) from University of Aveiro funded by the Portuguese Foundation for Science and Technology (FCT) through projects DOI: 10.54499/ UIDB/04106/2020 and DOI: 10.54499/UIDP/04106/2020. Additionally, JEFD acknowledges PhD contract DOI: 10.54499/UI/BD/152576/2022 from FCT. JEFD and MM acknowledge research project [PID2021- 122154NB-I00], Ortogonalidad y Aproximación con Aplicaciones en Machine Learning y Teoría de la Probabilidad funded by MICIU/AEI/10.13039/501100011033 and by “ERDF A Way of making Europe”.

### v2

@@ 5.2. Bidiagonal factorisation using the explicit expression for the type I polynomials
$\alpha_{(r+1)(rm+k)+i}=\frac{(a_{k+i+1}-a_{i}+m)(b+rm+k+1)\prod\limits_{j=1}^{r-1}\left(a_{i+j}+b+rm+k\right)}{\prod\limits_{j=1}^{r+1}\left(a_{k+i+j}+b+(r+1)m+k\right)},$ (5.21)
for the nontrivial entries of the bidiagonal matrices appearing in the factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ of the Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials on the step-line. It is a straightforward computation to check that (5.20)-(5.21) is equivalent to (5.8).
When $r=2$, the coefficients in (5.8) and (5.20)-(5.21) coincide with the ones found in [4, § 8.1].
It is a direct consequence of (5.8) or (5.20)-(5.21) that $\alpha_{n}\to\frac{r^{r}}{(r+1)^{r+1}}$ as $n\to\infty$. Therefore, recalling (2.10), the coefficients of the recurrence relation (2.7) satisfied by the type II Jacobi-Piñeiro polynomials on the step-line have the same limit as in [17, Eq. 8.17]: $\gamma_{n}^{[k]}\to\binom{r+1}{k+1}\left(\frac{r^{r}}{(r+1)^{r+1}}\right)^{k+1}$ as $n\to\infty$.
@@ 5.3. Multiple Laguerre of first kind as a limiting case
For a positive integer $r$ and parameters $a_{1},\cdots,a_{r}>0$ such that $a_{i}-a_{j}\not\in\mathbb{Z}$ whenever $i\neq j$, we consider the measures $\left(\nu_{1},\cdots,\nu_{r}\right)$ supported on the interval $(0,+\infty)$ with densities $\mathrm{d}\nu_{j}(x)=\mathrm{e}^{-x}\,x^{a_{j}-1}\mathrm{d}x$ for all $1\leq j\leq r$, whose moments are
$\int_{0}^{+\infty}x^{n}\mathrm{d}\nu_{j}(x)=\int_{0}^{+\infty}\mathrm{e}^{-x}\,x^{a_{j}+n-1}\,\mathrm{d}x=\Gamma(a_{j}+n)\quad\text{for all }n\in\mathbb{N}\text{ and }1\leq j\leq r.$ (5.22)
@@
The Hessenberg recurrence matrix of the multiple Laguerre polynomials of the first kind on the step-line, multiple orthogonal with respect to $(\nu_{1},\cdots,\nu_{r})$ given by (5.22), admits the factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$, as a product of the bidiagonal matrices in (1.2) with nontrivial entries $\left(\hat{\alpha}_{n}\right)_{n\in\mathbb{N}}$ such that $\hat{\alpha}_{n}=\lim\limits_{b\to\infty}b\,\alpha_{n}$ for all $n\in\mathbb{N}$, where $\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ are the nontrivial entries in the bidiagonal decomposition of the recurrence matrix of the Jacobi-Piñeiro polynomials given in (5.8) and (5.20)-(5.21). Therefore,
$\hat{\alpha}_{n}=a_{j(n)}-a_{i(n)}+\left\lfloor\dfrac{n}{r}\right\rfloor-\left\lfloor\dfrac{n}{r+1}\right\rfloor\quad\text{for all }n\in\mathbb{N},$ (5.26)
with $a_{0}=0$, $i(n)=[n]_{r+1}$, and $j(n)=[n]_{r}+1$, where $n]_{k}$ denotes the remainder of the division of $n$ by $k$, or, alternatively,
$\hat{\alpha}_{(r+1)(rm+k)+i}=\begin{cases}\hfil a_{k+i+1}-a_{i}+m&\text{if }k+i\leq r,\\ a_{k+i+1-r}-a_{i}+m&\text{if }k+i\geq r+1,\end{cases}$ (5.27)
for all $m\in\mathbb{N}$, $0\leq k\leq r-1$, and $0\leq i\leq r$.
It is a direct consequence of either (5.26) or (5.27) that $\hat{\alpha}_{n}\sim\frac{n}{r(r+1)}$ as $n\to\infty$. Therefore, recalling (2.10), the recurrence coefficients $\left(\hat{\gamma}_{n}^{[k]}\right)_{n\in\mathbb{N}}^{0\leq k\leq r}$ for the type II multiple Laguerre polynomials of the first kind on the step-line satisfy the asymptotic behaviour $\hat{\gamma}_{n}^{[k]}\sim\binom{r+1}{k+1}\left(\frac{n}{r}\right)^{k+1}$ as $n\to\infty$.
Note that, alternatively to taking the limit in (5.26), we could apply the algorithms developed in Sections 3 and 4 directly to the multiple Laguerre polynomials of first kind. For the method from Section 3, we could input the formulas for the type I multiple Laguerre polynomials of the first kind with respect to an arbitrary number of weights obtained in [6, Thm. 3] in (3.18)-(3.19) to obtain the coefficients $\left(\hat{\alpha}_{n}\right)_{n\in\mathbb{N}}$. However, it is much easier to simply take the limit in (5.26).
Using the algorithm from Section 4, the nontrivial entries $\left(\hat{\alpha}_{n}\right)_{n\in\mathbb{N}}$ of the bidiagonal decomposition of the recurrence matrix of the multiple Laguerre polynomials of first kind are the coefficients of the recurrence relation
@@
This relation is satisfied by the sequence $\left(\hat{g}_{n}(t)\right)_{n\in\mathbb{N}}$ obtained by taking the limit $\hat{g}_{n}(t)=\lim\limits_{b\to\infty}g_{n}(bt)$ for all $n\in\mathbb{N}$, where $\left(g_{n}(t)\right)_{k\in\mathbb{N}}$ is the sequence defined by (5.5). The computation of this limit is connected to contour integral representations for type I Jacobi-Piñeiro and multiple Laguerre polynomials and will be considered in forthcoming work.
@@ Acknowledgements
AB acknowledges financial support by the Centre for Mathematics of the University of Coimbra (CMUC, https://doi.org/10.54499/UID/00324/2025) under the Portuguese Foundation for Science and Technology (FCT), Grants UID/00324/2025 and UID/PRR/00324/2025. JEFD, AF, and HL acknowledge Center for Research and Development in Mathematics and Applications (CIDMA) from University of Aveiro funded by the Portuguese Foundation for Science and Technology (FCT) through grants UID/4106/2025 and UID/PRR/4106/2025. Additionally, JEFD acknowledges PhD contract DOI: 10.54499/UI/BD/152576/2022 from FCT. JEFD and MM acknowledge research project [PID2021- 122154NB-I00], Ortogonalidad y Aproximación con Aplicaciones en Machine Learning y Teoría de la Probabilidad funded by MICIU/AEI/ 10.13039/501100011033 and by “ERDF A Way of making Europe”.